OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 Think i did home work rite Can anyone check 4 me?

OpenStudy (anonymous):

#10 (3,O) Radius is 3 I did (x-3)^2+(y-0)^2=3^2

OpenStudy (anonymous):

you can just type: (x-3)^2+y^2=3^2

OpenStudy (anonymous):

Ok but our teach is picky lol

OpenStudy (anonymous):

well im sure he can accept that a number minus 0 is the number lol

OpenStudy (anonymous):

Yay i can acctualy do this lesson WOOOOOO A+ f me im usaly horibel!!

OpenStudy (anonymous):

*for

OpenStudy (anonymous):

a lot of times you have to complete the squares to get the standard equation of a circle, so you should know how to do that